At 9:00 a.m. a truck leaves the truck yard and travels west at a rate of 35 mi/hr two hours later a second truck leaves the same route traveling at 70 mi/hr When will the second truck catch up to the first
Answer by mananth(16946)

At 9:00 a.m. a truck leaves the truck yard and travels west at a rate of 35 mi/hr two hours later a second truck leaves the same route traveling at 70 mi/hr When will the second truck catch up to the first
..
Truck A 9.00 AM speed = 35 mph
Truck B 11.00 AM speed = 70mph
...
when Truck B starts Truck A in 2 hours has traveled = 35*2 = 70 miles
..
Both are travelling in same direction.
catch up speed = 70-35 = 35 mph
...
catch up distance / catch up speed = time
70/35 = 2hours
..
Truck B will catch up at 11.00 AM
...
m.ananth@hotmail.ca